40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16845     Entry time: Wed May 11 15:49:42 2022
Author: JC 
Type: Update 
Category: OPLEV Tables 
Subject: Green Beam OPLEV Alignment 

[Paco, JC]

Paco and I began aligning the Green Beam in the BS Oplev Table. while aligning the GRN-TRX, the initial beam was entering the table a bit low. To fix this, Paco went into the chamber and correcting the pitch with the steering mirror. The GRN-TRX is now set up, both the PD and Camera. Paco is continuing to work on the GRN-TRY and will update later on today. 

In the morning, I will update this post with photos of the new arrangement of the BS OPLEV Table.


Update Wed May 11 16:54:49 2022

[Paco]

GRY is now better mode matched to the YARM and is on the edge of locking, but it more work is needed to improve the alignment. The key difference this time with respect to previous attempts was to scan the two lenses on translation stages along the green injection path. This improved the GTRY level by a factor of 2.5, and I know it can be further improved. Anyways, the locked HOMs are nicely centered on the GTRY PD, so we are likely done with the in-vac GTRY GTRX alignment.


Update Wed May 12 10:59:22 2022

[JC]

The GTRX PD is now set up and connected. The camera have been set to an angle because the cable to connect it is too thick for the camera to maintain its original position along the side. 

 

Attachment 1: IMG_0770.jpeg  2.384 MB  Uploaded Thu May 12 11:58:40 2022  | Hide | Hide all
IMG_0770.jpeg
ELOG V3.1.3-